OpenStudy (anonymous):

2 to the zero power plus 2 to the neg 4 power. (simplify)

OpenStudy (anonymous):

17/16

OpenStudy (anonymous):

\[2^0+2^{-4} = 1+0.0002 = 1.0002 \]

OpenStudy (anonymous):

2^0=1 2^-4=1/16 1+1/16 make 1=16/16 16/16+1/16=17/16